# On some Sobolev and Pólya-Szegö type inequalities with weights and applications

### v3

@@ Abstract.
(P) $\begin{cases}-\Delta_{x}u-|x|^{2\alpha}\dfrac{\partial^{2}u}{\partial y^{2}}=f(x,y,u),&\textrm{in }\Omega,\\ u=0,&\textrm{on }\partial\Omega,\end{cases}$
@@ 1. Introduction
(1.1) $\begin{cases}-\dfrac{\partial^{2}u}{\partial x^{2}}-|x|^{2\alpha}\dfrac{\partial^{2}u}{\partial y^{2}}=f(x,y,u),&\textrm{in }\Omega,\\ u=0,&\textrm{on }\partial\Omega,\end{cases}$
@@
where $B_{s_{j}}:=\left\{(x_{1},x_{2},y)\in\mathbb{R}^{3}_{s_{j}},\textrm{ }\dfrac{|x|^{2\alpha+2}}{(\alpha+1)^{2}}+y^{2}<1\right\}$. Here, the notations $P_{2,\alpha}$, $P_{2,\alpha,j}$, and $|\cdot|_{2,\alpha}$ respectively denote weighted areas and weighted volume, which will be precisely given in the Definitions 1-3.
@@
(1.3) $\Bigg(\int\limits_{\mathbb{R}^{3}}|x|^{2\alpha}|u|^{6}dx_{1}dx_{2}dy\Bigg)^{1/6}\leq C^{-1}_{2\alpha,6}\Bigg(\int\limits_{\mathbb{R}^{3}}|\nabla_{G}u|^{2}dx_{1}dx_{2}dy\Bigg)^{1/2},$
@@
The other objective of this paper, achieved in Section 4, is to establish the existence and nonexistence results for the problem (P). In the case where the function $f$ is given by
@@
Assuming $f$ has the form given in (1.4). Let $\Omega$ be $G_{\alpha}$ – star-shaped with respect to the origin and $p>5$ (the definition of this type of domains will be given in Definition 7). Then, the problem (P) has no nontrivial solution $u\in\mathcal{S}_{2}(\overline{\Omega})$. Here, the function space $\mathcal{S}_{2}(\overline{\Omega})$ will be defined in the beginning of Subsection 4.1.
@@
    $\int_{\Omega}|\varphi(x_{1},x_{2},y)|dx_{1}dx_{2}dy<\infty,$
@@
Suppose that $f$ satisfies (A1)-(A5). Then, the problem (P) has a nontrivial weak solution (see Definition 11 for the definition of weak solutions).
@@ 2. A Pólya-Szegö type inequality
$\partial_{s_{j}}E:=\partial E\cap\mathbb{R}^{3}_{s_{j}},$
@@
$|x|:=\sqrt{x_{1}^{2}+x_{2}^{2}},\textrm{ for each }(x_{1},x_{2})\in\mathbb{R}^{2}.$
@@
$\dfrac{P_{2,\alpha}(E)^{3/2}}{|E|_{2,\alpha}}\textrm{ and }\dfrac{P_{2,\alpha,j}(E)^{3/2}}{|E|_{2,\alpha}},$
@@ 3. Sobolev-type inequalities with weights
(3.2) $\int\limits_{0}^{\infty}r^{m-1}|\phi^{\prime}(r)|^{p}dr<\infty,\textrm{ }\phi(r)\to 0\textrm{ when }r\to\infty.$
@@
$\dfrac{\big(\int_{0}^{\infty}r^{m-1}|\phi^{\prime}(r)|^{p}\big)^{1/p}}{\big(\int_{0}^{\infty}r^{m-1}|\phi(r)|^{q}\big)^{1/q}}\geq D_{p,q,m}$
@@ 4. Applications to a class of degenerate elliptic equations
In this section, we are mainly concerned with the existence and nonexistence of nontrivial solutions to the problem (P). Notice that our methods are similar to the ones that have been used in [3], and thus many details will be omitted.
@@ 4.1. Nonexistence result
In this subsection, we will derive the nonexistence of nontrivial solutions to the problem (P) for $f(x_{1},x_{2},y,\xi)=|x|^{2\alpha}|\xi|^{p-1}\xi$, $p\geq 1$. To this end, let
@@
A function $u(x_{1},x_{2},y)\in\mathcal{S}_{2}(\overline{\Omega})$ is said to be a solution to the problem (P) if
$\begin{cases}-\Delta_{x}u-|x|^{2\alpha}\dfrac{\partial^{2}u}{\partial y^{2}}=|x|^{2\alpha}|u|^{p-1}u,&\textrm{in }\Omega,\\ u=0,&\textrm{on }\partial\Omega.\end{cases}$
@@
Let $u(x_{1},x_{2},y)\in\mathcal{S}_{2}(\overline{\Omega})$ be a solution of (P). Denote by $\nu=(\nu_{1},\nu_{2},\nu_{3})$ the unit outward normal on $\partial\Omega$. Then we have
@@ 4.2. Existence result
In this subsection, we will give the existence result for the problem (P).
@@
$\|u\|_{L^{p}_{|x|^{2\alpha}}(\Omega)}:=\bigg(\int\limits_{\Omega}|x|^{2\alpha}|u|^{p}dx_{1}dx_{2}dy\bigg)^{1/p},$
@@
Now we are able to study the existence result for the problem (P).
@@
A function $u\in\mathcal{S}^{2}_{1,0}(\Omega)$ is called a weak solution of the problem (P) if the following identity
@@
In the following we define the energy functional $\Phi:\mathcal{S}^{2}_{1,0}(\Omega)\to\mathbb{R}$ associated with the problem (P) by letting
@@
$\langle\Phi^{\prime}(u),v\rangle=\int_{\Omega}\nabla_{G}u\cdot\nabla_{G}vdx_{1}dx_{2}dy-\int_{\Omega}f(x,y,u)vdx_{1}dx_{2}dy,$
for all $v\in\mathcal{S}^{2}_{1,0}(\Omega)$. Therefore, the weak solutions of the problem (P) are critical points of the functional $\Phi$.
We need to use the Mountain Pass Lemma to study the existence of critical points of the functional $\Phi$. First, let us recall the notion of the $(C)_{c}$ condition, which is an important notion in the statement of this lemma.
@@
We will need the following version of the Mountain Pass Lemma (see also [5, 6]).
@@
Acknowledgements
The second author of this paper is a member of a VAST project coded NVCC01.02/26-27.

### v4

@@ Abstract.
(P) $\begin{cases}-\Delta_{x}u-|x|^{2\alpha}\dfrac{\partial^{2}u}{\partial y^{2}}=f(x,y,u)&\textrm{in }\Omega,\\ u=0&\textrm{on }\partial\Omega,\end{cases}$
@@ 1. Introduction
(1.1) $\begin{cases}-\dfrac{\partial^{2}u}{\partial x^{2}}-|x|^{2\alpha}\dfrac{\partial^{2}u}{\partial y^{2}}=f(x,y,u)&\textrm{in }\Omega,\\ u=0&\textrm{on }\partial\Omega,\end{cases}$
@@
where $B_{s_{j}}:=\left\{(x_{1},x_{2},y)\in\mathbb{R}^{3}_{s_{j}},\textrm{ }\dfrac{|x|^{2\alpha+2}}{(\alpha+1)^{2}}+y^{2}<1\right\}$. Here, the notations $P_{2,\alpha}$, $P_{2,\alpha,j}$, and $|\cdot|_{2,\alpha}$ respectively denote weighted areas and weighted volume, which will be precisely given in Definitions 1-3.
@@
(1.3) $\Bigg(\int\limits_{\mathbb{R}^{3}}|x|^{2\alpha}|u|^{6}dx_{1}dx_{2}dy\Bigg)^{1/6}\leq C^{-1}_{2\alpha,6}\Bigg(\int\limits_{\mathbb{R}^{3}}|\nabla_{G}u|^{2}dx_{1}dx_{2}dy\Bigg)^{1/2}$
@@
The other objective of this paper, achieved in Section 4, is to establish the existence and nonexistence results for the problem (LABEL:P). In the case where the function $f$ is given by
@@
Assuming $f$ has the form given in (1.4). Let $\Omega$ be $G_{\alpha}$ – star-shaped with respect to the origin and $p>5$ (the definition of this type of domains will be given in Definition 7). Then, the problem (LABEL:P) has no nontrivial solution $u\in\mathcal{S}_{2}(\overline{\Omega})$. Here, the function space $\mathcal{S}_{2}(\overline{\Omega})$ will be defined in the beginning of Subsection 4.1.
@@
    $\int_{\Omega}|\varphi(x_{1},x_{2},y)|dx_{1}dx_{2}dy<\infty$
@@
Suppose that $f$ satisfies (A1)-(A5). Then, the problem (LABEL:P) has a nontrivial weak solution (see Definition 11 for the definition of weak solutions).
@@ 2. A Pólya-Szegö type inequality
$\partial_{s_{j}}E:=\partial E\cap\mathbb{R}^{3}_{s_{j}}$
@@
$|x|:=\sqrt{x_{1}^{2}+x_{2}^{2}}\textrm{ for each }(x_{1},x_{2})\in\mathbb{R}^{2}.$
@@
$\dfrac{P_{2,\alpha}(E)^{3/2}}{|E|_{2,\alpha}}\textrm{ and }\dfrac{P_{2,\alpha,j}(E)^{3/2}}{|E|_{2,\alpha}}$
@@ 3. Sobolev-type inequalities with weights
$\int\limits_{0}^{\infty}r^{m-1}|\phi^{\prime}(r)|^{p}dr<\infty,\textrm{ }\phi(r)\to 0\textrm{ when }r\to\infty.$
@@
(3.2) $\dfrac{\big(\int_{0}^{\infty}r^{m-1}|\phi^{\prime}(r)|^{p}\big)^{1/p}}{\big(\int_{0}^{\infty}r^{m-1}|\phi(r)|^{q}\big)^{1/q}}\geq D_{p,q,m}$
@@ 4. Applications to a class of degenerate elliptic equations
In this section, we are mainly concerned with the existence and nonexistence of nontrivial solutions to the problem (LABEL:P). Notice that our methods are similar to the ones that have been used in [3], and thus many details will be omitted.
@@ 4.1. Nonexistence result
In this subsection, we will derive the nonexistence of nontrivial solutions to the problem (LABEL:P) for $f(x_{1},x_{2},y,\xi)=|x|^{2\alpha}|\xi|^{p-1}\xi$, $p\geq 1$. To this end, let
@@
A function $u(x_{1},x_{2},y)\in\mathcal{S}_{2}(\overline{\Omega})$ is said to be a solution to the problem (LABEL:P) if
$\begin{cases}-\Delta_{x}u-|x|^{2\alpha}\dfrac{\partial^{2}u}{\partial y^{2}}=|x|^{2\alpha}|u|^{p-1}u&\textrm{in }\Omega,\\ u=0&\textrm{on }\partial\Omega.\end{cases}$
@@
Let $u(x_{1},x_{2},y)\in\mathcal{S}_{2}(\overline{\Omega})$ be a solution of (LABEL:P). Denote by $\nu=(\nu_{1},\nu_{2},\nu_{3})$ the unit outward normal on $\partial\Omega$. Then we have
@@ 4.2. Existence result
In this subsection, we will give the existence result for the problem (LABEL:P).
@@
$\|u\|_{L^{p}_{|x|^{2\alpha}}(\Omega)}:=\bigg(\int\limits_{\Omega}|x|^{2\alpha}|u|^{p}dx_{1}dx_{2}dy\bigg)^{1/p}$
@@
Now we are able to study the existence result for the problem (LABEL:P).
@@
A function $u\in\mathcal{S}^{2}_{1,0}(\Omega)$ is called a weak solution of the problem (LABEL:P) if the following identity
@@
In the following we define the energy functional $\Phi:\mathcal{S}^{2}_{1,0}(\Omega)\to\mathbb{R}$ associated with the problem (LABEL:P) by letting
@@
$\langle\Phi^{\prime}(u),v\rangle=\int_{\Omega}\nabla_{G}u\cdot\nabla_{G}vdx_{1}dx_{2}dy-\int_{\Omega}f(x,y,u)vdx_{1}dx_{2}dy$
for all $v\in\mathcal{S}^{2}_{1,0}(\Omega)$. Therefore, the weak solutions of the problem (LABEL:P) are critical points of the functional $\Phi$.
We need to use Mountain Pass Lemma to study the existence of critical points of the functional $\Phi$. First, let us recall the notion of the $(C)_{c}$ condition, which is an important notion in the statement of this lemma.
@@
We will need the following version of Mountain Pass Lemma (see also [5, 6]).